OpenStudy (anonymous):

Prove that (1-tan^2 A)/(1+tan^2 A) == cos2A

OpenStudy (anonymous):

\[{1-\tan^2A \over 1+\tan^2A}={1-\tan^2A \over \sec^2A}={1 \over \sec^2A}-{\sin^2A \over \cos^2A}\cdot \cos^2A\] \[=\cos^2A-\sin^2A=\cos(2A).\]

OpenStudy (y2o2):